what is percentage change in distance if the force of attraction between two point charges increases to 4 times keeping of charges constant?

Grade:12

1 Answers

Narendran
7 years ago
As,
F\alpha \frac{1}{r^{2}}.
 
r\alpha \frac{1}{\sqrt{F}}.
 
there for  reduces to half as F goes upto 4 times.Hence the percentage change is -50% (reduces).